# MAP-NBV: Multi-agent Prediction-guided Next-Best-View Planning for Active 3D Object Reconstruction

### v1

@@ I Introduction
Visual surveying and inspection with robots have been studied for a long time for a wide range of applications such as inspection of civil infrastructure [1, 2] and large vehicles [3, 4], precision agriculture [5], and digital mapping for real estate [6, 7]. The utilization of robots in these applications is highly advantageous as they can access hard-to-reach areas with greater ease and safety compared to situations with direct human involvement. Recent work on making robots autonomous for these tasks make their use more appealing. This work focuses on one such long-studied problem of 3D object reconstruction [8], where the objective is to digitally reconstruct the object of interest by combining observations from multiple vantage points. While it could be easier to achieve this in an indoor environment by carefully placing sensors around the object, the same can’t be achieved for the outdoors and open areas. For the latter, the sensor(s), must be moved around the object to capture information from different viewpoints. This can be realized with sensors such as cameras and LiDARs mounted on unmanned aerial vehicles (UAVs). A UAV with unlimited power supply capacity could capture infinite observations for an almost perfect reconstruction of the object, but the real-world limitation of battery capacity adds another dimension to the problem: achieving an accurate 3D reconstruction as fast as possible.
The trade-off between reconstruction accuracy and task duration in unknown environments is commonly addressed through Next-Best-View (NBV) planning, wherein a robot determines the optimal location for the next observation to maximize information gain. Numerous solutions have been proposed by the research community to tackle this problem, with a majority of them catering to single-agent systems [9]. However, deploying a team of robots instead of a single agent can enhance task efficiency multi-fold, while also offering additional benefits such as fault tolerance through redundancy. But the direct application of single-agent NBV methods to multi-agent systems does not translate well in terms of performance. This issue stems from the potential overlap between the individual observations. An efficient multi-agent NBV formulation requires coordination among robots to build a joint representation and minimize the overlap.
@@
In this work, we extend our previous work on prediction-driven single-agent NBV, Pred-NBV [10], to a team of robots for 3D reconstruction to bring the advantages of the prediction-guided approach to a multi-agent system. We call this multi-agent prediction-based next-best-view method MAP-NBV. Pred-NBV [10] uses a 3D point cloud prediction network along with a geometric NBV approach while also considering the control effort required for object reconstruction. An important feature of Pred-NBV is that it doesn’t require the partially observed point cloud to be centered at the full object center, an implicit assumption in many 3D reconstruction networks. Naively extending Pred-NBV to a team of robots would result in significant overlap as all the agents would move in the same direction to maximize individual information gain. This is inefficient as it would be more advantageous for the robots to move in different directions. MAP-NBV solves this issue by defining NBV measures over joint observation. We accomplish this by removing duplicate points in observations from multiple robots when calculating the information gain. Along with this, we account for the total control effort in our NBV objective, which results in efficient planning for the whole team.
@@
    We modify a single-agent baseline NBV algorithm based on [11] that uses frontier-based information gain, and extend its functionality to effectively operate in multi-agent settings.
@@
    We show that our method outperforms Pred-NBV [10], a single-agent prediction-based algorithm, by 22.75% and the multi-agent version of a traditional NBV baseline [11] by 15.63%.
@@ II Related Work
The use of robots for data acquisition purposes is an extensively studied topic for various domains. Their usage range from infrastructure inspection [12] and environment monitoring [13, 14] for real-world application to the real-world digitization for research datasets and simulations [6, 7, 15]. When the environment is unknown, active methods such as next-best-view (NBV) are used to construct an object model on the fly by capturing additional observations. A majority of the works on NBV planning use information-theoretic measures [9] for selection to account for uncertainty in observations [9, 16, 17]. The widely used frontier and tree-based exploration approaches also utilize uncertainty about the environment for guiding the robot motion [18, 19, 20, 21]. Some works devise geometric methods which make inferences about the exact shape of the object of interest and try to align the observations with the inferred model [22, 23, 24]. Prediction-based NBV approaches have emerged as another alternative in recent years, where a neural network takes the robot and/or the environment state as the input and NBV pose or velocity as the output [25, 26, 27, 10].
@@
A majority of the existing work on NBV is focused on single robot systems. The task performance can be enhanced by adding more robots to the systems, but directly extending single-robot NBV approaches to multi-robot systems may result in sub-optimal performance due to significant overlap in observations. This issue led to the development of exploration algorithms specifically for multi-robot systems [28, 29, 30] with information-theoretic measures for determining NBV. Some recent works on multi-robot systems have explored the utilization of predictions for improvement in task efficiency. Almadhoun et al. [31] designed a hybrid planner that switches between a classical NBV approach and a learning-based predictor for NBV selection but uses a partial model obtained by robot observations only. Wu et al. [32] use a point cloud prediction model for plants to use the predicted point cloud as an oracle leading to better results than the traditional approaches. This method uses entropy-based information gain measures for NBV and is designed for plant phenotyping with robotic arms. These methods do not consider the control effort required which is important for UAVs with energy constraints when deployed for observing large objects such as airplanes and ships. Also, these works employ information theoretic NBV approaches. We aim to explore a prediction-based approach for geometric NBV selection.
@@
In this work, we extend Pred-NBV [10] which also uses point cloud prediction and build a multi-robot NBV planner. The prediction on the point cloud makes the pipeline modular and interpretable and can be improved by improving individual modules. We select NBV based on information gain, as well as control effort, making our approach more grounded in real-world limitations.
@@ IV Proposed Approach
In this paper, we present Multi-Agent Pred-NBV (MAP-NBV), a model prediction-guided NBV approach for a team of robots. Figure 2 shows the overview of our process, which consists of two parts: (1)3D Model Prediction, where we combine the observations from all the robots to build a partial model of the object and use PoinTr-C [10], a 3D point cloud completion network, to predict the full shape of the objects, and (2) Multi-Agent NBV Algorithm, which uses the partial model and the predicted model to determine the NBV for the team, while trying to minimize the distance traveled. Our NBV solution performs a greedy selection over the candidate points to generate the trajectory, which also reduces the computation complexity. The following subsections provide further details of our approach.
[…]
@@
To start, the target object is segmented out from the rest of the environment in the captured RGB images for each UAV. This allows the algorithm to focus on only the target infrastructure as opposed to also including other obstacles. Then, each of these segmented images is aligned with the captured depth image per UAV to segment the target object out. Point clouds are then generated per each segmented depth image. This gives us a point cloud per each UAV that contains points belonging only to the target object. Assuming a centralized system, each segmented point cloud per UAV is transformed into a central reference frame and concatenated together into a singular point cloud. This point cloud represents the entire multi-agent system’s observations of the target object at the current timestamp. The point cloud concatenation can be replaced with a registration algorithm [33], but we use concatenation due to its ease of use. Lastly, this current timestamp’s point cloud is then concatenated with previous observations to get an up-to-date observation point cloud. This process is shown in Figure 3.
In order to get an approximation of the $\mathcal{\hat{V}}$ of the full model $\mathcal{V}$, we use PoinTr-C [10] a 3D point cloud completion network, developed by fine-tuning PoinTr [34] using curriculum learning over ShapeNet dataset [35]. Unlike PoinTr and similar point cloud completion networks, PoinTr-C doesn’t make implicit assumptions about the knowledge of the center of the full model by fine-tuning over rotationally and translationally perturbed point clouds. Relaxing this assumption makes PoinTr-C more suitable for inputs from an unknown environment than PoinTr. The 3D point cloud of the object obtained as the union of the observed surface points goes as input to PoinTr-C and it predicts the full object point cloud $\mathcal{\hat{V}}$.
PoinTr-C was trained over isolated point clouds and therefore requires object point clouds to be isolated from the scene. This can be realized with the help of distance-based filters and state-of-the-art segmentation networks[36] without any fine-tuning. An example of an input point cloud and a predicted point cloud is shown in Figure 4.
@@ IV-B Next-Best View Planner
The number of new points varies in each iteration since the robots observe more of the surface of the object as they move to new locations. While PoinTr-C predicts the point cloud for the whole object, the robots can observe only the surface points. Hence, before counting the number of new points, we apply hidden point removal [37] to the predicted point cloud. We represent this relationship between the number of points observed and the trajectories traversed till time $t$ by $I(\{\bar{\xi}_{t})$, where $\bar{\xi}_{t}=\{\xi_{r_{1}},\xi_{r_{2}},...,\xi_{r_{n}}\}_{t}$ represents the set of trajectories for all the robots till time $t$. To balance the information gain and control effort, we use a hyperparameter $\tau$ which is kept fixed throughout an episode. The robots select the candidate to pose set which results in at least $\tau\%$ of the total possible information gain over all candidate poses. Thus, we formulate our multi-agent NBV objective as follows.
@@ V-A Setup
We extend the setup in Pred-NBV [10] to work in a multi-agent setting. Similarly, we use Robot Operating System (ROS) Melodic and AirSim [39] on Ubuntu 18.04 for our simulation experiments. Multiple UAVs are spawned into the AirSim environment. We equipped each of the UAVs with a depth camera and an RGB camera. Each UAV published a segmented image using AirSim’s built-in segmentation. We adapted the depth segmentation package from Pred-NBV to work with multiple UAVs. We then converted these segmented depth images into 3D point clouds. For our collision-free point-to-point planner, we use the MoveIt [40] software package implementing the work done by Köse [41].
@@ V-C Comparison with Single-agent Baseline
We compared the performance of MAP-NBV with a single-agent prediction-based NBV planner called Pred-NBV [10]. MAP-NBV is an extension of Pred-NBV designed for multi-agent scenarios. However, in single-agent cases, both algorithms function identically. In MAP-NBV, UAVs are spawned close together, ensuring that the initial environment information is virtually the same as in the single-agent Pred-NBV case. Consequently, the initial points observed and the initial shape completion predictions for both algorithms are highly similar. This means that MAP-NBV and Pred-NBV select their initial NBVs using nearly identical information.
@@
TABLE I: Points observed by MAP-NBV and Pred-NBV [10] after first iteration of algorithm.
@@ V-D Comparison with Multi-agent Baseline
We also compared the performance of MAP-NBV with a modified baseline NBV method [11] designed for multi-agent use. The baseline method employs frontiers to select the next-best views. Frontiers are points located at the edge of the observed space near unknown areas. We utilized the same modifications described in Pred-NBV [10]. Specifically, we used our segmented point cloud to choose frontiers near the target object. To ensure that the UAVs always face the target object, the orientation of all poses selected by the baseline aligns with the center of the observed target object point clouds.
@@
TABLE II: Points observed by MAP-NBV and the multi-agent baseline NBV method [11] for all models in AirSim upon algorithm termination.
@@
Figure 7: Comparison between MAP-NBV and the Multi-agent baseline NBV algorithm [11] for the Church Tower.

### v3

@@ I Introduction
Visual surveying and inspection with robots have been studied for a long time for a wide range of applications such as inspection of civil infrastructure [shanthakumar2018view, dhamiGATSBI] and large vehicles [kim_2009_IROS, ropek_2021], precision agriculture [dhami2020crop], and digital mapping for real estate [46965, ramakrishnan2021hm3d]. The utilization of robots in these applications is highly advantageous as they can access hard-to-reach areas with greater ease and safety compared to situations with direct human involvement. Recent work on making robots autonomous for these tasks make their use more appealing. This work focuses on one such long-studied problem of 3D object reconstruction [bajcsy2018revisiting], where the objective is to digitally reconstruct the object of interest by combining observations from multiple vantage points. While it could be easier to achieve this in an indoor environment by carefully placing sensors around the object, the same can’t be achieved for the outdoors and open areas. For the latter, the sensor(s), must be moved around the object to capture information from different viewpoints. This can be realized with sensors such as cameras and LiDARs mounted on unmanned aerial vehicles (UAVs). A UAV with unlimited power supply capacity could capture infinite observations for an almost perfect reconstruction of the object, but the real-world limitation of battery capacity adds another dimension to the problem: achieving an accurate 3D reconstruction as fast as possible.
The trade-off between reconstruction accuracy and task duration in unknown environments is commonly addressed through Next-Best-View (NBV) planning, wherein a robot determines the optimal location for the next observation to maximize information gain. Numerous solutions have been proposed by the research community to tackle this problem, with a majority of them catering to single-agent systems [delmerico2018comparison]. However, deploying a team of robots instead of a single agent can enhance task efficiency multi-fold, while also offering additional benefits such as fault tolerance through redundancy. But the direct application of single-agent NBV methods to multi-agent systems does not translate well in terms of performance. This issue stems from the potential overlap between the individual observations. An efficient multi-agent NBV formulation requires coordination among robots to build a joint representation and minimize the overlap.
@@
In this work, we extend our previous work on prediction-driven single-agent NBV, Pred-NBV [dhami2023prednbv], to a team of robots for 3D reconstruction to bring the advantages of the prediction-guided approach to a multi-agent system. We call this multi-agent prediction-based next-best-view method MAP-NBV. Pred-NBV [dhami2023prednbv] uses a 3D point cloud prediction network along with a geometric NBV approach while also considering the control effort required for object reconstruction. An important feature of Pred-NBV is that it doesn’t require the partially observed point cloud to be centered at the full object center, an implicit assumption in many 3D reconstruction networks. Naively extending Pred-NBV to a team of robots would result in significant overlap as all the agents would move in the same direction to maximize individual information gain. This is inefficient as it would be more advantageous for the robots to move in different directions. MAP-NBV solves this issue by defining NBV measures over joint observation. We accomplish this by removing duplicate points in observations from multiple robots when calculating the information gain. Along with this, we account for the total control effort in our NBV objective, which results in efficient planning for the whole team.
@@
    We modify a single-agent baseline NBV algorithm based on [aleotti2014global] that uses frontier-based information gain, and extend its functionality to effectively operate in multi-agent settings.
@@
    We show that our method outperforms Pred-NBV [dhami2023prednbv], a single-agent prediction-based algorithm, by 22.75% and the multi-agent version of a traditional NBV baseline [aleotti2014global] by 15.63%.
@@ II Related Work
The use of robots for data acquisition purposes is an extensively studied topic for various domains. Their usage range from infrastructure inspection [ozaslaninspection] and environment monitoring [dunbabin2012environmental, sung2019competitive] for real-world application to the real-world digitization for research datasets and simulations [46965, ramakrishnan2021hm3d, ammirato2017dataset]. When the environment is unknown, active methods such as next-best-view (NBV) are used to construct an object model on the fly by capturing additional observations. A majority of the works on NBV planning use information-theoretic measures [delmerico2018comparison] for selection to account for uncertainty in observations [delmerico2018comparison, kuipers1991robot, vasquez2014volumetric]. The widely used frontier and tree-based exploration approaches also utilize uncertainty about the environment for guiding the robot motion [yamauchi1997frontier, gonzalez2002navigation, adler2014autonomous, bircher2018receding]. Some works devise geometric methods which make inferences about the exact shape of the object of interest and try to align the observations with the inferred model [tarabanis1995survey, banta2000next, kriegel2013combining]. Prediction-based NBV approaches have emerged as another alternative in recent years, where a neural network takes the robot and/or the environment state as the input and NBV pose or velocity as the output [johns2016pairwise, mendoza2020supervised, zeng2020pc, dhami2023prednbv].
@@
A majority of the existing work on NBV is focused on single robot systems. The task performance can be enhanced by adding more robots to the systems, but directly extending single-robot NBV approaches to multi-robot systems may result in sub-optimal performance due to significant overlap in observations. This issue led to the development of exploration algorithms specifically for multi-robot systems [burgard2005coordinated, amanatiadis2013multi, hardouin2020next] with information-theoretic measures for determining NBV. Some recent works on multi-robot systems have explored the utilization of predictions for improvement in task efficiency. Almadhoun et al. [almadhoun2021multi] designed a hybrid planner that switches between a classical NBV approach and a learning-based predictor for NBV selection but uses a partial model obtained by robot observations only. Wu et al. [wu2019plant] use a point cloud prediction model for plants to use the predicted point cloud as an oracle leading to better results than the traditional approaches. This method uses entropy-based information gain measures for NBV and is designed for plant phenotyping with robotic arms. These methods do not consider the control effort required which is important for UAVs with energy constraints when deployed for observing large objects such as airplanes and ships. Also, these works employ information theoretic NBV approaches. We aim to explore a prediction-based approach for geometric NBV selection.
@@
In this work, we extend Pred-NBV [dhami2023prednbv] which also uses point cloud prediction and build a multi-robot NBV planner. The prediction on the point cloud makes the pipeline modular and interpretable and can be improved by improving individual modules. We select NBV based on information gain, as well as control effort, making our approach more grounded in real-world limitations.
@@ IV Proposed Approach
In this paper, we present Multi-Agent Pred-NBV (MAP-NBV), a model prediction-guided NBV approach for a team of robots. Figure 2 shows the overview of our process, which consists of two parts: (1)3D Model Prediction, where we combine the observations from all the robots to build a partial model of the object and use PoinTr-C [dhami2023prednbv], a 3D point cloud completion network, to predict the full shape of the objects, and (2) Multi-Agent NBV Algorithm, which uses the partial model and the predicted model to determine the NBV for the team, while trying to minimize the distance traveled. Our NBV solution performs a greedy selection over the candidate points to generate the trajectory, which also reduces the computation complexity. The following subsections provide further details of our approach.
(a) Input
(b) PoinTr-C Prediction
@@
To start, the target object is segmented out from the rest of the environment in the captured RGB images for each UAV. This allows the algorithm to focus on only the target infrastructure as opposed to also including other obstacles. Then, each of these segmented images is aligned with the captured depth image per UAV to segment the target object out. Point clouds are then generated per each segmented depth image. This gives us a point cloud per each UAV that contains points belonging only to the target object. Assuming a centralized system, each segmented point cloud per UAV is transformed into a central reference frame and concatenated together into a singular point cloud. This point cloud represents the entire multi-agent system’s observations of the target object at the current timestamp. The point cloud concatenation can be replaced with a registration algorithm [huang2021comprehensive], but we use concatenation due to its ease of use. Lastly, this current timestamp’s point cloud is then concatenated with previous observations to get an up-to-date observation point cloud. This process is shown in Figure 3.
In order to get an approximation of the $\mathcal{\hat{V}}$ of the full model $\mathcal{V}$, we use PoinTr-C [dhami2023prednbv] a 3D point cloud completion network, developed by fine-tuning PoinTr [yu2021pointr] using curriculum learning over ShapeNet dataset [shapenet2015]. Unlike PoinTr and similar point cloud completion networks, PoinTr-C doesn’t make implicit assumptions about the knowledge of the center of the full model by fine-tuning over rotationally and translationally perturbed point clouds. Relaxing this assumption makes PoinTr-C more suitable for inputs from an unknown environment than PoinTr. The 3D point cloud of the object obtained as the union of the observed surface points goes as input to PoinTr-C and it predicts the full object point cloud $\mathcal{\hat{V}}$.
PoinTr-C was trained over isolated point clouds and therefore requires object point clouds to be isolated from the scene. This can be realized with the help of distance-based filters and state-of-the-art segmentation networks[kirillov2023segment] without any fine-tuning. An example of an input point cloud and a predicted point cloud is shown in Figure 4.
@@ IV-B Next-Best View Planner
The number of new points varies in each iteration since the robots observe more of the surface of the object as they move to new locations. While PoinTr-C predicts the point cloud for the whole object, the robots can observe only the surface points. Hence, before counting the number of new points, we apply hidden point removal [katz2007direct] to the predicted point cloud. We represent this relationship between the number of points observed and the trajectories traversed till time $t$ by $I(\{\bar{\xi}_{t})$, where $\bar{\xi}_{t}=\{\xi_{r_{1}},\xi_{r_{2}},...,\xi_{r_{n}}\}_{t}$ represents the set of trajectories for all the robots till time $t$. To balance the information gain and control effort, we use a hyperparameter $\tau$ which is kept fixed throughout an episode. The robots select the candidate to pose set which results in at least $\tau\%$ of the total possible information gain over all candidate poses. Thus, we formulate our multi-agent NBV objective as follows.
@@ V-A Setup
We extend the setup in Pred-NBV [dhami2023prednbv] to work in a multi-agent setting. Similarly, we use Robot Operating System (ROS) Melodic and AirSim [airsim2017fsr] on Ubuntu 18.04 for our simulation experiments. Multiple UAVs are spawned into the AirSim environment. We equipped each of the UAVs with a depth camera and an RGB camera. Each UAV published a segmented image using AirSim’s built-in segmentation. We adapted the depth segmentation package from Pred-NBV to work with multiple UAVs. We then converted these segmented depth images into 3D point clouds. For our collision-free point-to-point planner, we use the MoveIt [coleman2014reducing] software package implementing the work done by Köse [tahsinko86:online].
@@ V-C Comparison with Single-agent Baseline
We compared the performance of MAP-NBV with a single-agent prediction-based NBV planner called Pred-NBV [dhami2023prednbv]. MAP-NBV is an extension of Pred-NBV designed for multi-agent scenarios. However, in single-agent cases, both algorithms function identically. In MAP-NBV, UAVs are spawned close together, ensuring that the initial environment information is virtually the same as in the single-agent Pred-NBV case. Consequently, the initial points observed and the initial shape completion predictions for both algorithms are highly similar. This means that MAP-NBV and Pred-NBV select their initial NBVs using nearly identical information.
@@
TABLE I: Points observed by MAP-NBV and Pred-NBV [dhami2023prednbv] after first iteration of algorithm.
@@ V-D Comparison with Multi-agent Baseline
We also compared the performance of MAP-NBV with a modified baseline NBV method [aleotti2014global] designed for multi-agent use. The baseline method employs frontiers to select the next-best views. Frontiers are points located at the edge of the observed space near unknown areas. We utilized the same modifications described in Pred-NBV [dhami2023prednbv]. Specifically, we used our segmented point cloud to choose frontiers near the target object. To ensure that the UAVs always face the target object, the orientation of all poses selected by the baseline aligns with the center of the observed target object point clouds.
@@
TABLE II: Points observed by MAP-NBV and the multi-agent baseline NBV method [aleotti2014global] for all models in AirSim upon algorithm termination.
@@
Figure 7: Comparison between MAP-NBV and the Multi-agent baseline NBV algorithm [aleotti2014global] for the Church Tower.
@@ VI Conclusion
VII Appendix
O: observation
P: planning
SO(1)-IP: Share observation 1-hop, Independented planing
SO(c)-SP: Shared observation (centralized), Independent planning
Optimal: SO(c)-SP(c)-Optimal SeqGreedy: SO(c)-SP(c)-Greedy DecGreedyv1: SO(1)-IP DecGreedyv2: SO(n)-SP(n)-Greedy Independent: IO-IP
Algorithm 1 Optimal NBV Assignment Algorithm
1:Inputs: Initial positions $x_{i,0},\forall i\in R$; Stopping threshold $\tau$
2:Output: Point cloud of the object under inspection
3:$novelty$ $\leftarrow 0$; $t\leftarrow 1$; $O_{comb,0}\leftarrow\phi$; #$\#points_{0}\leftarrow 1$
4:while $novelty\leq\tau$ do
5:     $O_{t}\leftarrow\bigcup_{i\in R}$ getObservedPointCloud($x_{i}$)
6:     $\hat{O}\leftarrow$ completePointCloud($O_{t}$)
7:     $C\leftarrow$ generateCandidiatePoses($\hat{O}$)
8:     for all possible assignments $\hat{X}=\{0\text{:}c_{i},...,R\text{:}c_{j}\}$ do
9:         $p_{comb,\hat{X}}=\bigcup_{i\in R}$ getVisiblePoints($i\text{:}c$)
10:     end for
11:     $X_{NBV}\leftarrow\underset{\hat{X}}{\operatorname*{argmax}}\,p_{comb,\hat{X}}$
12:     $X\leftarrow X_{NBV}$
13:     #$\#points_{t}\leftarrow|p_{comb,X_{NBV}}|$
14:     $novelty\leftarrow\frac{\#points_{t}}{\#points_{t-1}}$
15:     $t\leftarrow t+1$
16:end while
Algorithm 2 Sequential Greedy Assignment Algorithm
1:Inputs: Initial positions $x_{i,0},\forall i\in R$; Stopping threshold $\tau$
2:Output: Point cloud of the object under inspection
3:$novelty$ $\leftarrow 0$; $t\leftarrow 1$; $O_{comb,0}\leftarrow\phi$; #$\#points_{0}\leftarrow 1$
4:while $novelty\leq\tau$ do
5:     $O_{t}\leftarrow\bigcup_{i\in R}$ getObservedPointCloud($x_{i}$)
6:     $\hat{O}\leftarrow$ completePointCloud($O_{t}$)
7:     $C\leftarrow$ generateCandidiatePoses($\hat{O}$)
8:     $points_{seen}\leftarrow\phi$
9:     $X_{NBV}\leftarrow\phi$
10:     for $i\leftarrow$ 1 to R do
11:         $x_{NBV,i}\leftarrow\underset{c\in C\setminus X_{NBV}}{\operatorname*{argmax}}\,$ getVisiblePoints($c$) $\setminus points_{seen}$
12:         $x_{i}\leftarrow x_{NBV,i}$
13:         $X_{NBV}\leftarrow X_{NBV}\bigcup x_{i}$
14:         $points_{seen}\leftarrow points_{seen}\bigcup$ getVisiblePoints($x_{i}$)
15:     end for
16:     $X\leftarrow X_{NBV}$
17:     #$\#points_{t}\leftarrow$ getVisiblePoints($X_{NBV}$)
18:     $novelty\leftarrow\frac{\#points_{t}}{\#points_{t-1}}$
19:     $t\leftarrow t+1$
20:end while
Algorithm 3 Decentralized Greedy Assignment Algorithm (1-hop communication, run on each robot)
1:Inputs: Initial position $x$; Stopping threshold $\tau$
2:Output: Next Best View positions $x_{NBV}$
3:$novelty$ $\leftarrow 0$; $t\leftarrow 1$; $O_{comb,0}\leftarrow\phi$; #$\#points_{0}\leftarrow 1$
4:while $novelty\leq\tau$ do
5:     $O_{t}\leftarrow\bigcup_{i\in Neighbors}$ getObservedPointCloud($x_{i}$)
6:     $\hat{O}\leftarrow$ completePointCloud($O_{t}$)
7:     $C\leftarrow$ generateCandidiatePoses($\hat{O}$)
8:     $points_{seen}\leftarrow\phi$
9:     $x_{NBV}\leftarrow\underset{c\in C}{\operatorname*{argmax}}\,$ (getVisiblePoints($c$) $\setminus points_{seen}$)
10:     $x\leftarrow x_{NBV}$
11:     $x\leftarrow X_{NBV}$
12:     $points_{seen}\leftarrow points_{seen}\bigcup$ getVisiblePoints($x_{i}$)
13:     #$\#points_{t}\leftarrow$ getVisiblePoints($X_{NBV}$)
14:     $novelty\leftarrow\frac{\#points_{t}}{\#points_{t-1}}$
15:     $t\leftarrow t+1$
16:end while
Algorithm 4 Decentralized Greedy Assignment Algorithm (multi-hop communication)
Runs Sequential greedy on each subgraph